OpenStudy (lina17):

how do you do this problem y over -6 +5=9 by step by step

OpenStudy (lina17):

\[\frac{ y }{ -6} +5=9\]

OpenStudy (lina17):

the answers are A.24 B.-24 C.78 D.-78

OpenStudy (sleepyjess):

Hello! Welcome to OpenStudy! To do this, we need to isolate y on the left side of the equal sign. To start that, first we need to subtract 5 from both sides. Can you show me that?

OpenStudy (lina17):

so i subtract 5 from 9

OpenStudy (sleepyjess):

Yep :)

OpenStudy (lina17):

so it would be like this \[\frac{ y }{ -6 }=4\]

OpenStudy (sleepyjess):

\(\dfrac y{-6} + 5 - 5 = 9 - 5\) + 5 - 5 = 0 9 - 5 = 4 That gives us \(\dfrac y{-6} = 4\)

OpenStudy (sleepyjess):

Good! You're ahead of me already haha :)

OpenStudy (sleepyjess):

Now, we need to cancel out -6. We subtracted 5 to cancel out addition, so we divide by -6 to cancel the multiplication. Does that make sense?

OpenStudy (anonymous):

Bruh, did you read what you wrote?

OpenStudy (sleepyjess):

Whoops, my mind was a little backwards, we multiply by -6 to cancel the division

OpenStudy (lina17):

that does not make sense

OpenStudy (lina17):

so y =24

OpenStudy (sleepyjess):

Close, don't forget there's a negative sign :)

OpenStudy (lina17):

i mean -24

OpenStudy (sleepyjess):

Yep :3

OpenStudy (lina17):

lol thanks

OpenStudy (lina17):

can you help me out on one more please

OpenStudy (sleepyjess):

Sure!

OpenStudy (lina17):

its 4(x+1)+8=24

OpenStudy (sleepyjess):

Okay, for this one, let's distribute the 4 first

OpenStudy (lina17):

so you subtract 4 from both sides ?

OpenStudy (sleepyjess):

Not really... Focus on the 4(x + 1). We need to multiply 4*x and 4*1

OpenStudy (lina17):

oh okay

OpenStudy (lina17):

so it would be 4x+5

OpenStudy (sleepyjess):

So after doing that, we have 4x + 4 + 8 = 24. This will be easier if we combine the 4 + 8, giving 4x + 12 = 24.

OpenStudy (lina17):

its 4x+5+8=24

OpenStudy (sleepyjess):

What is 4 * 1?

OpenStudy (lina17):

oh lol sorry

OpenStudy (sleepyjess):

Anything multiplied by 1 is itself :)

OpenStudy (sleepyjess):

It's fine

OpenStudy (lina17):

so it would be 4x+12=24

OpenStudy (sleepyjess):

Yep :) Now this is just like the last one. How do we cancel the +12 on the left side?

OpenStudy (lina17):

so we subtract the 4 on both sides

OpenStudy (sleepyjess):

We're not quite there yet, we need to cancel the 12 first

OpenStudy (lina17):

how do i do that

OpenStudy (lina17):

sorry i keep forgetting because my daughter keeps distracting me

OpenStudy (sleepyjess):

Oh no! I'm so sorry! I went to go get my math book and got distracted. Basically, whenever we need to cancel something, we want to do that opposite of what is originally being done. If we're adding, subtract that amount; if we're subtracting, add that amount; if we're multiplying, divide by that number; if we're dividing, multiply by that number.

OpenStudy (lina17):

ok

OpenStudy (lina17):

so how do i cancel out the 12

OpenStudy (lina17):

@sleepyjess

OpenStudy (xxkesha):